The winery has a corresponding tasting room at an offsite location. In 2015, they ordered approximately 1200 bottles of wine approximately once a month, which was based on the average annual sales demand data for 2015 (R=14708 per year). If their ordering cost is K = $400 per shipment and their annual holding cost is h = $0.50 per bottle, would you recommend that they change their order quantity in 2016, or should they stick with ordering approximately 1200 bottles of wine each month in 2016 as they have been doing for 2015? Please provide a detailed discussion and analysis, addressing any concerns you have with continued use of their current order quantity.
Answer #1

Annual Demand = D = 14708/year

Ordering Cost = Co = $400

Holding Cost = Cc = $0.50/year

Economic Order Quantity Q* = √(2DCo/Cc) = √(2*14708*400/0.50) = 4851.06 or 4851

Annual Inventory Holding cost = average inventory * unit carrying cost = (Q*/2)*Cc = (4851/2)*0.50 = $1212.75

Annual Order cost = Number of orders * cost/order = (D/Q*) Co = (14708/4851)*400 = $1212.78

Total Annual Cost = Annual Inventory Holding Cost + Annual Order Cost = 1212.75 + 1212.78 = $2425.53

If order Quantity Q = 1200,

Total Annual Cost = (Q*/2)*Cc + (D/Q*) Co = (1200/2)*0.50 + (14708/1200)*400 = $5202.67

The total cost is higher when 1200 bottles are ordered by 5202.67 - 2425.53 = $2777.14

Hence, the order quantity should be changed to 4851